OpenStudy (anonymous):

Given f(x) = x – 10, what is f(–7)? A. –17 B. –3 C. –10 D. –7

OpenStudy (solomonzelman):

plug in -7 for x

OpenStudy (sidsiddhartha):

f(x)=x-10 f(-7)=-7-10 just pluf in -7 in place of x

OpenStudy (anonymous):

@sidsiddhartha -7 - 10?

OpenStudy (sidsiddhartha):

-17

OpenStudy (anonymous):

Thanks!